# Surrogate Modeling of 3D Rayleigh-Bénard Convection with Equivariant Autoencoders

### v1

@@ 3.2 Rayleigh-Bénard convection
3.2.1 Symmetries of the 3D Rayleigh-Bénard convection system
@@
The PDE we consider is a modification of the Navier-Stokes equations (NSE) with an additional temperature-dependent buoyancy term in the vertical direction. It is well-known that the 3D NSE are equivariant under actions of the symmetry group E(3) (Olver, 1993; Wang et al., 2021). Consequently, our system inherits the translational equivariance and $O(2)$ equivariance in the two horizontal directions, as only the vertical component of the momentum equation is altered. In addition, the translational symmetry in the vertical direction is broken by the fixed-temperature boundary conditions at the bottom and top. Locally, however, and sufficiently far away from the walls, equivariance should approximately hold, which we will exploit in our architecture. A more detailed discussion—including a proof—can be found in Appendix A.1.
4 Methods
4.1 3D steerable convolution on E(2)
As discussed in Section 3.2.1, the 3D Rayleigh-Bénard convection is $E(2)$-equivariant in the horizontal plane, which we enforce by steerable convolutions, while introducing height-dependent kernels that adapt to the varying dynamics at different heights.
4.1.1 Steerable convolution
The system state $s$ consists of both scalar temperature and vector-valued velocity fields. Under transformation of these fields by $\tau g\in E(2)$, the temperature field $T:\mathbb{R}^{3}\rightarrow\mathbb{R}$ transforms as $T(x)\mapsto 1\cdot T(g^{-1}(x-\tau))$, whereas the velocity-field $u:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ transforms as $u(x)\mapsto g\cdot u(g^{-1}(x-\tau))$. Note that the velocity vectors are themselves transformed via $g$ to preserve their orientation as the field is transformed Cohen and Welling (2017); Weiler and Cesa (2019).
Thus, a three-dimensional $O(2)$-steerable kernel can be constructed as a stack of 2D $O(2)$-steerable kernels. For these, Weiler and Cesa (2019) have solved the steerability constraint as well as for important subgroups such as $C_{4}$ and $D_{4}$. This result has been applied in our implementation to efficiently design height-dependent kernels utilizing the PyTorch-based library escnn. [^1] https://github.com/QUVA-Lab/escnn For computational reasons, we restrict our implementation to the subgroup $H=(\mathbb{Z}^{2},+)\rtimes D_{4}<E(2)$, as this optimally corresponds to our data on a rectangular grid. Thus, we will from now on consider discretized kernels and feature fields on $\mathbb{Z}^{3}$. The case of continuous rotations according to $O(2)$ would require interpolation between grid points (see, e.g., Esteves et al. (2018a)). In our experiments, this resulted in inferior performance compared to the grid-consistent 90-degree-rotations and flips.
Within our framework, we use various types of feature fields. Both the input to the AE-encoder and the output of the decoder are composed of the scalar field $T$ and the vector field $u$. All intermediate representations, however, make use of regular feature fields, which transform under the regular representation by permuting the channels. Steerable convolutions between regular feature fields are equivalent to regular group convolutions (Cohen and Welling, 2016), which apply the same kernel in every orientation $g\in G$, resulting in the $|G|$-dimensional regular feature fields.
4.1.2 3D convolutions with height-dependent kernels
Height-dependent features—such as temperature or velocity patterns—play a critical role in modeling the system. As a result, applying the same kernel across all heights—as would be the case for a regular 3D CNN—is insufficient to capture the system’s vertical dynamics. To address this limitation, we modify the conventional 3D convolutions by learning height-dependent steerable kernels. While this ensures horizontal parameter sharing, we allow the convolution operation to adapt to the distinct features at each height, ensuring that the vertical structure is captured effectively. For computing the output feature map’s value at position $x=(x_{1},x_{2},x_{3})\in\mathbb{Z}^{3}$, the input $f:\mathbb{Z}^{3}\rightarrow\mathbb{R}^{C_{\mathsf{in}}}$ is convolved with the height-dependent kernel $\psi_{x_{3}}:\mathbb{Z}^{3}\rightarrow\mathbb{R}^{C_{\mathsf{out}}\times C_{\mathsf{in}}}$ via $[f*\psi](x)=\sum_{y\in\mathbb{Z}^{3}}\psi_{x_{3}}(x-y)f(y)$.
Local vertical parameter sharing. Although the Rayleigh-Bénard system does not exhibit global vertical translation equivariance, it approximately maintains this property within a local neighborhood (see Section 3.2.1). This suggests that features at a given height $x_{3}$ are locally correlated with features at vertical positions within a $12k+$ 1-sized neighborhood $\mathcal{N}({x}_{3})=\{{x}_{3}-k,\ldots,{x}_{3}+k\}$. This approximate local equivariance in the vertical direction can be exploited by choosing a smaller number of channels in each layer, but then applying these learned kernels across the local neighborhood $\mathcal{N}$, thereby again increasing the number of output channels, cf. Fig. 3 for a sketch. This parameter reduction leads to a more efficient and scalable model, rendering it suitable for simulating larger-scale systems. As described in Appendix A.2, 3D convolutions with height-dependent kernels and local vertical parameter sharing can be efficiently implemented by wrapping 2D (steerable) convolutions.
4.2 Equivariant autoencoder
The decoder applies a series of convolutions, followed by activation functions. Pooling operations are replaced by upsampling. To preserve continuity while upsampling, we employ trilinear interpolation. Throughout the entire framework (CAE + LSTM), padding is applied to the convolutional layers to ensure that the spatial dimensions of the data are only altered by pooling and upsampling operations. For the horizontal dimensions, circular padding is used to match the periodic boundary conditions of the Rayleigh-Bénard system, while vertical dimensions use zero padding. Note that without padding, the decoder would require transposed convolutions (Dumoulin and Visin, 2018) to reverse the dimensional changes induced by convolutions without padding.
4.3 Equivariant LSTM
Long short-term memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) extend conventional recurrent neural networks (RNNs) for processing sequential data by an additional cell state $c_{t}$. While the hidden state $h_{t}$ stores currently relevant information for predicting the next time step, the cell state $c_{t}$ captures long-term information over long sequences. The content of $c_{t}$ is regulated by the input gate $i_{t}$, which controls which information is added, and the forget gate $f_{t}$, which controls how much information is discarded. The output gate $o_{t}$ determines which information is passed from $h_{t}$ to $c_{t}$. To preserve the spatial structure of the latent space $z$, we use convolutional LSTMs (Shi et al., 2015) that replace the fully connected layers of standard LSTMs with, in our case, equivariant convolutions and also introduce equivariant convolutions in peephole connections. The equations are defined as follows:
$h_{t}$ $=o_{t}\odot\tanh\left(c_{t}\right)$
$z_{t+1}=z_{t}+h_{t}*\psi+b.$
$(z_{t},c_{t-1},h_{t-1})\mapsto(z_{t+1},c_{t},h_{t}),$ (6)
It follows from three facts, all due to Weiler and Cesa (2019): Firstly, convolutions are equivariant in the sense that $\rho(g)\ell*\psi=\rho(g)(\ell*\psi)$ for an arbitrary feature map $\ell$. Secondly, the action is linear, i.e., $\rho(g)\ell+\rho(g)\ell^{\prime}=\rho(g)(\ell+\ell^{\prime})$. And thirdly, it satisfies $\sigma(\rho(g)\ell)=\rho(g)\sigma(\ell)$ whenever $\sigma$ is a pointwise function. Since all feature maps are transformed jointly, the Hadamard product is equivariant too, i.e., $\rho(g)\ell\odot\rho(g)\ell^{\prime}=\rho(g)(\ell\odot\ell^{\prime})$.
End-to-end equivariance of the entire architecture then simply follows from the observation that all CAE as well as all LSTM layers are equivariant w.r.t. the same symmetry group.
5 Experiments
We compare the performance of our end-to-end equivariant framework to a similar-in-size baseline model using standard, non-steerable 3D convolutions, with a particular focus on the autoencoder.
Datasets.
We generated a dataset of 100 randomly initialized 3D Rayleigh-Bénard convection simulations with $Ra=2500$ and $Pr=0.7$, standardized to zero mean and unit standard deviation. The dataset was split into 60 training, 20 validation, and 20 test simulations, each with 400 snapshots in the time interval $t\in[100,300]$ at a step size of 0.5 and a spatial resolution of $N=48\times 48\times 32$. For long-term forecasting evaluation, we also created an additional dataset of 20 simulations, each containing 1800 snapshots over $t\in[100,1000]$.
Architecture.
The CAE encoder and decoder each consist of six convolutional layers with a kernel size of three for the last encoder and first decoder layer, and five for the other layers, applying ELU nonlinearities pointwise. [^2] All hidden layers in our model transform under the regular representation, which preserves equivariance since it commutes with pointwise nonlinearities (Weiler and Cesa, 2019). Pooling and upsampling are applied after the encoder and decoder’s first, third, and fifth layers, respectively. The latent space has dimensions $M=6\times 6\times 4$ and $C_{\mathsf{latent}}=32$ channels, resulting in a compression to 1.56% of the original size. Channels roughly double after each pooling operation in the encoder, with the decoder reversing this. Over our experiments, we vary the channel sizes. For a fair comparison, our main models (both standard and equivariant) have approximately the same number of 3.6M parameters. The decoder-only LSTM with one layer of convolutional LSTM cells and an additional convolutional layer for output prediction uses a kernel size of 3. The number of channels is chosen such that the LSTM has approximately 3.7M parameters.
Training.
CAE training is performed in the standard self-supervised manner on a set of snapshots, where the desired decoder-output is the original input. The LSTM was then trained on the latent representations to predict the 50 latent states following the provided sequence of 25 states, with the loss being computed on the latent space. This two-step training approach effectively separates encoding and forecasting, avoiding significant performance drops. Mixing them leads to poor results, as the LSTM output (the latent representation) is autoregressively fed back into the model without intermediate decoding and encoding. Consequently, end-to-end fine-tuning resulted in suboptimal forecasting performance. For more training details, cf. A.3.
5.1 Results
Autoencoder.
Fig. 4 gives an overview of our experiments to compare the equivariant and baseline architectures with respect to parameter efficiency, data efficiency, compression capabilties and scalability to more complex dynamics, i.e., larger $Ra$. Equivariance leads to a significant improvement in reconstruction accuracy, with the RMSE decreasing by 42%, from 0.069 $\pm$ 0.00072 to 0.04 $\pm$ 0.00093, where the mean and standard deviation were computed across three models with randomly initialized weights. A comparison to $C_{4}$-equivariant convolutions shows an RMSE of 0.046 $\pm$ 0.00164, showing that both rotations and reflections are relevant for the improved performance.
Figure 4: CAE performances. Results are shown with varying numbers of parameters, amounts of training samples, compression ratios, and Rayleigh numbers $Ra$. Each plot shows a variation over one of the parameters while keeping the others fixed. The main model has 3.6 million parameters ($3^{\mathsf{rd}}$ point), 24,000 training samples ($4^{\mathsf{th}}$ point), 1.56% latent size ($3^{\mathsf{rd}}$ point) and $Ra=2500$ ($1^{\mathsf{st}}$ point).
  1. (i)
    Parameter efficiency: The equivariant model with only 900,000 parameters outperformed the baseline model with 10.8 million parameters, meaning that we obtain an improvement by more than one order of magnitude.
  2. (ii)
    Sample efficiency: The $D_{4}$ model achieved the same performance as the baseline when trained on just one-eighth of the training samples, representing a significant reduction in the amount of data required for effective learning, making it particularly valuable in data-limited scenarios or when simulations/experiments are expensive.
  3. (iii)
  4. (iv)
Figure 5: Forecasting RMSE. For each of three randomly initialized models, the median forecasting RMSE was computed over a time horizon of 500. We present the average of these medians.
Long-term forecasting.
We next focus on the evaluation of the long-term forecasting capabilities of our end-to-end equivariant architecture. The model is provided with an input sequence of 50 snapshots and then predicts the subsequent 500 future states in an autoregressive manner, i.e., using its own predictions as input for the next time step.
Fig. 5 shows the median RMSE over time for both the equivariant and baseline model, which was averaged over three separately trained models. The equivariant model consistently outperforms the baseline by a near-constant margin of approximately 0.05 RMSE across the entire forecast horizon. This indicates that most of the performance gains stem from improved latent representation learned by the equivariant autoencoder, which provides a more stable and informative basis for prediction.
6 Conclusion
We have developed an end-to-end equivariant CAE plus LSTM architecture for efficient surrogate modeling of 3D Rayleigh-Bénard convection, which includes horizontally $E(2)$-equivariant kernels that are vertically stacked. Additional local sharing in the vertical direction allows us to increase the number of channels without requiring additional parameters, which further addes to the computational efficiency. We have demonstrated significant advantages in terms of the accuracy and both data and parameter efficiency.
6.1 Limitations and outlook
There are still several limitations that we have not yet discussed or addressed, but believe to be important and fruitful for future research
  • Instead of the sequential training approach, an end-to-end training could be superior, since the latent representation is tailored to the dynamics; since this proved to be challenging for LSTMs, an alternative could be linear models based on the Koopman operator (Harder et al., 2024a), which has proven to be a powerful tool for learning dynamical systems, also in combination with autoencoders (Otto and Rowley, 2019).
  • In the control setting, equivariant models are very helpful to improve the efficiency, for instance of world models in the context of reinforcement learning; surrogate modeling for RL of PDEs was studied in, e.g., Werner and Peitz (2024), but a combination with equivariant RL as proposed in van der Pol et al. (2020) has not been investigated.
Code
Acknowledgments
Appendix A Appendix
A.1 Symmetries in the Rayleigh-Bénard system
@@
The symmetries described in Section 3.2.1 can be formalized as equivariance with respect to three-dimensional rigid transformations $R(x)=Ax+b$, where $A$ leaves the vertical direction invariant, i.e., it has a $2\times 2$ orthogonal block in the top-left corner and acts like the identity in the third (= the vertical) dimension. Together with function composition as the operation, the set of these rigid transformations yields a group $(G,\circ)$.
@@
$\nabla(R\cdot w)$ $=R\cdot\nabla w,$ (for $w$ a scalar or vector field) (9)
@@
A.2 Implementation details
@@
To parallelize this across all $H_{\mathsf{out}}$ heights, we stack the receptive fields along the channel dimension, producing a tensor of shape $B\times(H_{\mathsf{out}}\cdot h\cdot C_{\mathsf{in}})\times N_{1}\times N_{2}$, and then apply a grouped 2D convolution, dividing the channels into $H_{\mathsf{out}}$ groups, each corresponding to a separate height (not to be confused with group-equivariant convolutions [Cohen and Welling, 2016]). After processing, the output tensor is reshaped back to $B\times C_{\mathsf{out}}\times N_{1}^{\prime}\times N_{2}^{\prime}\times H_{\mathsf{out}}$.
@@
A.3 Training details and computing resources
Both models were trained using the Adam optimizer [Kingma and Ba, 2017] with an MSE loss, a batch size of 64, a learning rate of $1\times 10^{-3}$, and learning rate decay. Training was stopped after convergence of the validation loss. Dropout, $D_{4}$ data augmentation, and batch normalization (which was applied only to the autoencoder) were used. The LSTM was trained using backpropagation through time and scheduled sampling [Bengio et al., 2015] for gradual transition from teacher forcing [^3] Teacher forcing uses the ground truth as the LSTM input during training, instead of its previous output. to autoregressive predictions.

### v2

@@ A.1 Boussinesq approximation and numerical solution
A.2 Symmetries in the Rayleigh-Bénard system
@@
The PDE we consider is a modification of the Navier-Stokes equations (NSE) with an additional temperature-dependent buoyancy term in the vertical direction. It is well-known that the 3D NSE are equivariant under actions of the symmetry group E(3) (Olver, 1993; Wang et al., 2021). Consequently, our system inherits the translational equivariance and $O(2)$ equivariance in the two horizontal directions, as only the vertical component of the momentum equation is altered. In addition, the translational symmetry in the vertical direction is broken by the fixed-temperature boundary conditions at the bottom and top. Locally, however, and sufficiently far away from the walls, equivariance should approximately hold, which we will exploit in our architecture.
@@
The symmetries can be formalized as equivariance with respect to three-dimensional rigid transformations $R(x)=Ax+b$, where $A$ leaves the vertical direction invariant, i.e., it has a $2\times 2$ orthogonal block in the top-left corner and acts like the identity in the third (= the vertical) dimension. Together with function composition as the operation, the set of these rigid transformations yields a group $(G,\circ)$.
@@
A.3 Additional systems with broken E(3) symmetry
Even though the paper is concerned with Rayleigh Bénard convection only, there exists a multitude of systems where our architecture can be of use. Examples include, but are not limited to:
  • systems with directed forces, such as
  • systems with symmetry-breaking flow structures such as
    • the flow through pipelines, where the symmetry is broken along the flow direction, leading to O(2) symmetry in radial direction and a shift equivariance along the transport direction.
    • jet flows such as the exhaustion gas coming out of a jet engine, where the symmetry is broken along the jet direction. As the jet increases in diameter in downstream direction, the remaining symmetry would be O(2) around the jet center axis.
Appendix B Architectural and implementation details of the surrogate and baselines
B.1 LSTM equations and equivariance
Our convolutional LSTM architecture is defined as follows:
$h_{t}$ $=o_{t}\odot\tanh\left(c_{t}\right)$
$z_{t+1}=z_{t}+h_{t}*\psi+b.$
$(z_{t},c_{t-1},h_{t-1})\mapsto(z_{t+1},c_{t},h_{t}),$ (16)
It follows from three facts, all due to Weiler & Cesa (2019): Firstly, convolutions are equivariant in the sense that $\rho(g)\ell*\psi=\rho(g)(\ell*\psi)$ for an arbitrary feature map $\ell$. Secondly, the action is linear, i.e., $\rho(g)\ell+\rho(g)\ell^{\prime}=\rho(g)(\ell+\ell^{\prime})$. And thirdly, it satisfies $\sigma(\rho(g)\ell)=\rho(g)\sigma(\ell)$ whenever $\sigma$ is a pointwise function. Since all feature maps are transformed jointly, the Hadamard product is equivariant too, i.e., $\rho(g)\ell\odot\rho(g)\ell^{\prime}=\rho(g)(\ell\odot\ell^{\prime})$.
B.2 Implementation details
@@
To parallelize this across all $H_{\mathsf{out}}$ heights, we stack the receptive fields along the channel dimension, producing a tensor of shape $B\times(H_{\mathsf{out}}\cdot h\cdot C_{\mathsf{in}})\times N_{1}\times N_{2}$, and then apply a grouped 2D convolution, dividing the channels into $H_{\mathsf{out}}$ groups, each corresponding to a separate height (not to be confused with group-equivariant convolutions (Cohen & Welling, 2016)). After processing, the output tensor is reshaped back to $B\times C_{\mathsf{out}}\times N_{1}^{\prime}\times N_{2}^{\prime}\times H_{\mathsf{out}}$.
@@
B.3 Baseline architectures
In addition to our proposed equivariant surrogate, we implemented two widely used baseline architectures, 3D U-Nets and 3D Fourier Neural Operators (FNOs). Their configurations are detailed below.
3D U-Net.
We implement a 3D U-Net with four downsampling and four upsampling stages and skip connections between encoder and decoder. Each block uses two 3D convolutions with a kernel size of three and ReLU nonlinearities. Downsampling is performed via max pooling; upsampling uses trilinear interpolation. The final layer is a $1{\times}1{\times}1$ convolution to the target channels. We set the base width to $C_{0}{=}28$ hidden channels, with channels being doubled/halved after each downsampling/upsampling step, respectively, yielding 7.7M parameters.
3D Fourier Neural Operator (FNO).
We use the neuralop (Kovachki et al., 2023; Kossaifi et al., 2024) 3D FNO implementation with 16 Fourier modes, 4 spectral layers, 20 hidden channels, GELU nonlinearities, and 7.4M parameters. Each FNO layer applies a spectral convolution in Fourier space (truncating to the specified modes) plus a pointwise linear transform in physical space. We do not use dropout and apply an $L_{2}$ weight decay of $1{\times}10^{-4}$ during training.
Appendix C Training and methodological considerations
C.1 Training details and computing resources
Both models were trained using the Adam optimizer (Kingma & Ba, 2017) with an MSE loss, a batch size of 64, a learning rate of $1\times 10^{-3}$, and learning rate decay. Training was stopped after convergence of the validation loss. Dropout, $D_{4}$ data augmentation, and batch normalization (which was applied only to the autoencoder) were used. The LSTM was trained using backpropagation through time and scheduled sampling (Bengio et al., 2015) for gradual transition from teacher forcing [^3] Teacher forcing uses the ground truth as the LSTM input during training, instead of its previous output. to autoregressive predictions.
@@ C.1 Training details and computing resources
Training times at different autoregressive horizons. To quantify the efficiency gap between our approach and the baselines, we also report average training times per epoch for the $D_{4}$-steerable LSTM, FNO, and U-Net. As shown in Table 1, the computational cost of FNOs and U-Nets grows dramatically with longer horizons, whereas our model remains efficient because forecasting is performed entirely within the latent space. This motivates training our models with 50 autoregressive steps, while FNO and U-Net baselines were restricted to 5 steps (see Section 5).
Table 1: Training times per epoch (in seconds) for different architectures. Compared to FNOs and U-Nets, our latent-space $D_{4}$-steerable model trains an order of magnitude faster when using longer autoregressive horizons, since forecasting is performed directly in latent space. This highlights its scalability for efficient long-term forecasting.
Training steps $D_{4}$-steerable LSTM (ours) [s] FNO [s] U-Net [s]
5 $582$ $916$ $1091$
50 $1310$ $8876$ 12 93912\,939
C.2 Discussion on two-step training approach
In this work, we adopted a two-step training strategy in which the autoencoder (CAE) and the LSTM are trained separately. While end-to-end training may appear more natural, our experiments showed that this approach offers several important advantages. Below, we provide a detailed discussion of this design choice, including the challenges we observed with end-to-end training and possible future improvements.
Challenges of end-to-end training
  • Mixing of compression and forecasting. When trained jointly, the LSTM can attempt to correct errors made by the CAE. This leads to unstable behavior in long-horizon autoregressive forecasts, as forecasting and reconstruction tasks interfere with each other.
  • High computational cost. End-to-end training requires backpropagation through the encoder, LSTM, and decoder over time, making it both computationally expensive and memory-intensive.
One possible mitigation is to decode the LSTM’s prediction, re-encode it, and feed it back into the LSTM for each autoregressive step. However, this introduces extreme inefficiency. On the other hand, separate training has multiple advantages.
Advantages of separate training
  • Clear separation of tasks. In end-to-end training, the model tends to integrate forecasting logic into the autoencoder, which is especially harmful for autoregressive forecasting in latent space over long horizons. By contrast, training the CAE and LSTM separately ensures that the CAE focuses purely on compression and reconstruction, while the LSTM is specialized for temporal dynamics.
  • Efficiency. Training the LSTM on pre-computed latent representations allows the forecasting loss to be computed directly in latent space, without backpropagating through the decoder and encoder. This reduced training time per epoch by approximately a factor of 20 (0.36 hours vs. 7.52 hours per epoch in our experiments), while also significantly lowering memory requirements.
  • Dynamical systems perspective. From a modeling point of view, the two-step approach aligns with the idea of first learning a low-dimensional manifold that represents the system dynamics (Connor & Rozell, 2020), and then training a forecasting model restricted to that manifold.
We believe that separate training could be enhanced by a careful end-to-end fine-tuning step. This would allow the latent representation to adapt to the forecasting task while retaining the benefits of efficient pre-training. Such an approach, however, requires a careful balancing of reconstruction and forecasting losses, as well as extensive hyperparameter tuning.
C.3 Advantages over classical PDE solvers
A key motivation for surrogate modeling is to overcome the large computational cost of classical PDE solvers, particularly for high-dimensional turbulent systems such as Rayleigh–Bénard convection. In this appendix, we provide additional measurements and discuss the advantages of our approach compared to numerical solvers.
Multi-query capability. Classical solvers must integrate each trajectory individually, which becomes prohibitively expensive in scenarios such as uncertainty quantification, optimization, or control. By contrast, the surrogate model can efficiently predict large batches in parallel, enabling applications that are infeasible with traditional PDE solvers.
Inference speed. We benchmarked our surrogate model against the highly optimized solver oceananigans.jl. For the Rayleigh number considered in this work, the surrogate already shows a modest speedup, while maintaining high accuracy. More importantly, the computational cost of classical solvers grows rapidly with increasing complexity, such that scaling towards more turbulent regimes will strongly favor machine learning approaches.
Table 2: Average inference times (500-step forecast) of our surrogate model ($D_{4}$-steerable CAE+LSTM) compared to the PDE solver oceananigans.jl. Reported times include only the forecasting computation (no memory or SSD I/O). While single-trajectory runtimes are comparable, the surrogate provides dramatic speedups in multi-query settings, where large batches can be simulated in parallel.
Scenario Ours [s] Solver [s] Speedup
Single trajectory, full sequence output 18.38 28.34 $1.5\times$
Single trajectory, only final state output 12.06 25.31 $2.1\times$
256 trajectories, full sequence output 257.12 ($\approx$1.00 / sim.) $28.3\times$
256 trajectories, only final state output 47.05 ($\approx$0.18 / sim.) $140.6\times$
Table 2 summarizes the average speedup factors for computation (CAE+LSTM vs. oceananigans.jl). The results highlight the clear advantage of surrogate models in multi-query settings, where large ensembles of trajectories can be simulated efficiently in parallel. In addition, our approach provides a substantial computational benefit when only the final state is required, since forecasting can be performed entirely in latent space and only the last state needs to be decoded.
Differentiability. Another strength of neural surrogates is their differentiability. Unlike most classical numerical solvers, which are either not differentiable (in particular high-performance solvers for very large simulations and commercial codes) or require the tedious implementation of adjoint equations, [^4] It should be noted that there have various attempts in the recent past to design fully differentiable solvers, in particular using the backpropagation functionalities included in modern ML packages such as PyTorch; cf., e.g., Holl & Thuerey (2024); Franz & Thuerey (2024); Winchenbach & Thuerey (2025). our model can be directly differentiated end-to-end. This enables gradient-based optimization, data assimilation, control, and parameter identification. We believe that the combination of scalability, efficiency, and differentiability makes surrogate models particularly attractive for downstream tasks in scientific machine learning.